A high school is planning to build a new field surrounded by a running track. The perimeter of the track is 500m.A rectangle field inside track with maximum area, what are the dimensions ?
Answer by Alan3354(69443)

Max area of a rectangle is a square (proof can be supplied).
Side = 500/4 = 125 m
Area = 125*125 = 15625 sq meters